Calculate wave number and frequency of blue radiation having wavelength 4800
A
 

Grade:11

3 Answers

Khimraj
5 years ago
-3)= 0.21x 108)/(3 x105Wave number =Frequency/c = (0.625x 105)/4800 = 0.625x 108Frequency =c/ Wavelength = (3 x10
Khimraj
5 years ago
-3)= 0.21x 108)/(3 x105Wave number =Frequency/c = (0.625x 105)/4800 = 0.625x 108Frequency =c/ Wavelength = (3 x10
Hope it clears.........................
Khimraj
5 years ago
Frequency =c/ Wavelength = (3 x108)/4800 = 0.625x 105
Wave number =Frequency/c = (0.625x 105)/(3 x108)= 0.21x 10-3